OpenStudy (anonymous):

Which is a recursive formula for the sequence 99.4, 0, –99.4, –198.8, where f(1) = 99.4

OpenStudy (anonymous):

f(n + 1) = f(n) + 99.4, n ≥ 1 f(n + 1) = f(n) – 99.4, n ≥ 1 f(n + 1) = 99.4f(n), n ≥ 1 f(n + 1) = –99.4f(n), n ≥ 1